OpenStudy (anonymous):

Please help me with my Algebra 2??? what is an equation of a parabola with the given vertex and focus vertex:(-2,5) focus:(-2,6) someone PLEASE help me :(

OpenStudy (anonymous):

do you remember the standard formula

jimthompson5910 (jim_thompson5910):

the distance from vertex to focus is 1 unit, so p = 1 the vertex is (-2,5), so (h,k) = (-2,5) ---> h = -2, k = 5 4p(y-k) = (x-h)^2 4*1(y-k) = (x-h)^2 4*1(y-5) = (x-(-2))^2 4*1(y-5) = (x+2)^2 y-5 = (1/4)(x+2)^2 I'll let you finish up

OpenStudy (anonymous):

I don't really know what to do

jimthompson5910 (jim_thompson5910):

you would solve for y and get it into vertex form or standard form (depending on what your teacher wants)

jimthompson5910 (jim_thompson5910):

vertex form is y = a(x-h)^2 + k

jimthompson5910 (jim_thompson5910):

standard form is y = ax^2 + bx + c

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

OpenStudy (anonymous):

Do I add 5 to both sides?

jimthompson5910 (jim_thompson5910):

yes that will get the equation into vertex form

jimthompson5910 (jim_thompson5910):

if necessary, you'll have to convert it to standard form

OpenStudy (anonymous):

I don't understand what to do after that

OpenStudy (anonymous):

As you can tell math is NOT my strong suit

OpenStudy (anonymous):

And I also have no idea how to find the focus

jimthompson5910 (jim_thompson5910):

the focus is given to you, in this case it's (-2,6)

jimthompson5910 (jim_thompson5910):

if you add 5 to both sides, you get y-5 = (1/4)(x+2)^2 y = (1/4)(x+2)^2+5 then you would expand out (x+2)^2, distribute the 1/4, then combine like terms

OpenStudy (anonymous):

y = (1/4)(x+2)^2+5? Is that the answer?

jimthompson5910 (jim_thompson5910):

that's the answer in vertex form

jimthompson5910 (jim_thompson5910):

if that's what they want, you can stop there

OpenStudy (anonymous):

So do I multiply by 4 to cancel the fraction?

jimthompson5910 (jim_thompson5910):

what do you get when you expand out (x+2)^2

OpenStudy (anonymous):

It just wants the equation of the parabola

OpenStudy (anonymous):

2x+4?

jimthompson5910 (jim_thompson5910):

(x+2)^2 = (x+2)(x+2) = ???

OpenStudy (anonymous):

Or x^2+4

jimthompson5910 (jim_thompson5910):

closer, but it's really x^2 + 2x + 4

jimthompson5910 (jim_thompson5910):

oops i meant x^2 + 4x + 4

jimthompson5910 (jim_thompson5910):

so y = (1/4)(x+2)^2+5 turns into y = (1/4)(x^2 + 4x + 4) + 5

OpenStudy (anonymous):

I kinda understand the math, but I don't know what the equation is? Do I use the vertex form?

jimthompson5910 (jim_thompson5910):

then you distribute the 1/4 through to get y = (1/4)(x^2 + 4x + 4) + 5 y = (1/4)x^2 + (1/4)*4x + (1/4)*4 + 5 y = (1/4)x^2 + x + 1 + 5 now combine like terms to get this y = (1/4)x^2 + x + 6

OpenStudy (anonymous):

So is that my equation?

OpenStudy (anonymous):

Thank you very very much for your help